# Uncovering the role of semantic and acoustic cues in normal and dichotic listening

### v1

@@ II-F Semantic Feature - Word2vec text embedding
The text transcription of the speech stimulus is provided in the dataset. This study used this text data to obtain features representing the semantics of the stimulus speech signal. Semantic vectors for content words were derived using the word2vec algorithm [19].
@@ II-G1 Acoustic Encoding
The EEG sub-network (top row of Figure 2) consists of a series of convolutional layers and LSTM layers. The convolutional layers implement -D and -D convolutions with and kernel sizes, respectively. Further, the -D CNN layers also introduce a stride of to further down-sample the feature maps. The speech sub-network (middle row in Figure 2) has a 2D convolutional layer of kernel size 16 and stride (1,3). For acoustic encoding, we utilize both the speech and EEG subnetworks in a parallel neural pipeline, as depicted in Figure 2, without any weight sharing.
The word boundary information available in the dataset is provided at the equivalent sampling rate (both EEG and audio representations at Hz). The audio and EEG feature maps are average pooled at the word level using the word boundary information. As a result, for a given sentence, the EEG and speech branches generate vector representations sampled at the word level. An LSTM layer models the inter-word context from these representations. This layer is included in both the stimulus (speech) and response (EEG) pathways. The last hidden state of the LSTM layer, of dimension , is used as the embedding for the stimulus/response, denoted as /, respectively.
We propose the Manhattan distance [39] between the stimulus and response embeddings. The similarity score is computed as,
The similarity score for the matched pair () and mismatched pair () are computed. The model, with a dropout factor of , is trained using a binary cross-entropy loss, with [, ] mapped to [, ] targets.
@@ II-G2 Semantic Encoding
To compare the encoding of text features with speech features, we used a text-EEG match mismatch classification task. We used the text-EEG model sub-network (bottom row of Figure 2) to learn the mapping between the w2v embeddings and the EEG response. The EEG sub-network resembles the acoustic encoding model, while the text sub-network consists of two-layers of LSTM architecture with w2v embeddings at word-level as input features. The convolution layers were not used as the word2vec (w2v) features do not contain 2-D information. The similarity scoring component is identical to the acoustic encoding model (Eq. 1).
@@ II-G3 Joint Encoding of Acoustics and Semantics
In addition to exploring the effect of acoustic and semantic features individually, we perform an investigation of the joint training of the MM model with both the modalities. The sub-networks are combined by defining a joint-loss function over the two modalities (detailed in Sec. II-I). For a given match segment, the EEG response corresponds to the attended audio/text features, while in the mis-matched example, the audio/text features are chosen from a different region (natural speech) or from the stream played in the other ear (dichotic listening).
@@ II-H Trial selection in Dichotic listening
As part of the dichotic listening task, participants were required to answer multiple-choice questions after each trial to assess their comprehension and, consequently, their attentiveness to the played stories [4]. They were asked questions each from both the stories. The scores obtained were normalized to a range of to . For all the experiments that we report in this work, we include trials in which the participants achieved a comprehension score of % or more for attended audio, while achieving a score of % or less for the un-attended audio. This choice of scores for attended / un-attended streams ensured that the “matched” / “mis-matched” EEG-audio/text pairs had a clear distinction. The selection is elicited in Figure 3. While this selection led to a removal of % of the total trials, the selection ensured a reliable label for the training and evaluation.
@@ II-H1 Training and Evaluation Setup
The experiments are run with a batch size of . The models are trained using Adam optimizer with a learning rate of and weight decay parameter of . The models are learned with a binary cross-entropy loss.
Overall, the natural speech dataset contains stimulus trials, while the dichotic listening dataset contains trials. A subset of trials from all subjects constituted the training set, while the remaining unseen trials of all subjects served as the test data. In each validation fold, we randomly chose trials to form the test data. This resulted in -fold cross-validation for the natural speech dataset and -fold cross-validation for the cocktail party dataset. The average classification accuracy across validation folds is reported in the results section.
@@ II-I Multimodal Loss Function and Joint training
Similarity | |||||||
measure | 0 | 0.2 | 0.4 | 0.5 | 0.6 | 0.8 | 1 |
82.00 | 82.12 | 82.87 | 83.76 | 82.92 | 79.46 | 77.45 | |
81.20 | 81.60 | 80.00 | 78.46 | 78.00 | 77.60 | 75.71 | |
77.76 | 79.00 | 78.72 | 78.12 | 77.57 | 76.97 | 74.61 |
A hyper parameter is introduced, which controls the contribution of the text and speech modalities in the joint training. We explore three loss function choices which combine the modalities,
Here, implies that only the semantic network (text-EEG) is considered, implies that only the acoustic network (speech-EEG) is considered while implies joint modeling (EEG-speech+text) with equal weight to speech and text modalities. During training, the speech, text and EEG data are loaded and the value of is randomly selected from . In this way, a single model is trained with all three combinations, of text alone (), speech alone () and speech-text multimodal (). During inference, the model can be input with EEG-text samples (), EEG-speech samples () as well EEG-speech+text (). This flexibility helps us understand the role of acoustics and semantics separately from the same model.
@@ III-A Choice of loss function
For dichotic listening condition, we experiment with different values during inference. As shown in Table I, it is observed that semantic encoding in EEG achieves a higher MM performance than the acoustic encoding network. Further, the similarity measure proved to be best performing similarity measure. The best MM classification results are achieved with . All the subsequent experiments use this configuration with similarity measure.
@@ III-B Impact of Acoustic and Semantic Cues
cells = c,
hlines,
vline2 = -,
hline1,4,6 = -0.08em,
hline5 = -dashed,
hline3 = -dashed,
colsep = 4pt
Listening
Condition & Speech
envelope
(=1) Text
word2vec
(=0) Multimodal
(Acoustic
+Semantic)
(=0.5)
Natural (Baseline) [12] 65.23 – –
Dichotic (Baseline) [33] 62.00 – –
Natural 93.63 93.24 93.38
Dichotic 77.45 82.00 83.76
Table II presents the results of the MM classification accuracy between speech stimuli and their corresponding EEG responses, focusing on different listening conditions. Compared to the MM classification of natural speech under normal listening conditions, the task of auditory attention decoding (AAD) in dichotic listening is more challenging.
For natural speech conditions, both semantic and acoustic features result in similar accuracy (Wilcoxon signed rank test, ). For dichotic listening, word2vec features result in significant improvement in accuracy compared to speech envelope features (Wilcoxon signed rank test, ). The scatter plot for prediction probabilites for dichotic listening condition and for a fold in validation is shown in Figure 4.
@@ III-B1 Importance of word boundary information
Word-level segmentation holds significant importance in speech perception during normal listening conditions [17]. To investigate the role of word segmentation in auditory attention decoding, we conducted a comparison of match-mismatch classification performance using the proposed model with and without word boundary information in the dichotic listening condition. The word boundary segmentation is used only for the EEG and acoustic features, while the word2vec features are inherently sampled at word-level.
As seen in Table III, word level segmentation plays an important role in auditory attention decoding. Word boundary information has impact on both semantic features (Wilcoxon rank-sum test, ) and envelope features (Wilcoxon rank-sum test, ). Word boundary information also shows a significant impact on the joint speech-text model.
column2-3 = c,
hlines,
hline3-4 = -dashed,
vline2-3 = -
Stimulus Feature & Without
word-boundary With
word-boundary
Envelope (=1) 67.09 77.45
word2vec (=0) 77.35 82.00
Multi-modal (=0.5) 76.69 83.76
@@ III-B2 Importance of accurate word boundaries
We conducted another ablation study to understand the impact of accurate word boundary information. In this setting, the model is fed with random word boundaries with varying granularity. Each sentence is segmented to a fixed number of regions, and their boundaries are chosen at random. The results are reported in Table IV. The performance does not significantly improve with acoustic feature (Wilcoxon rank-sum test, ), semantic feature (Wilcoxon rank-sum test, ), or with multi-modality (Wilcoxon rank-sum test, ) as we increase the number of random word boundaries. However, the use of accurate word-boundaries is significantly better in all cases over the ones with random word boundaries.
row1 = l,
cell11 = l,
cell22-6 = c,
cell32-6 = c,
cell42-6 = c,
hlines,
vline2 = -,
vline6 = -dashed,
hline3-4 = -dashed,
Stimulus
Feature & 2 3 4 5 True word
boundary
Envelope
(=1)
66.13 66.16 67.39 70.89 77.45
word2vec
(=0)
78.95 76.29 78.58 78.19 82.00
Multi-modal
(=0.5)
79.11 79.20 79.55 79.41 83.76
In the second set of ablation experiments, we provide accurate word boundary information but skip the word boundary information at every -th word. These results are reported in Table V. For example, ‘Skip-3’ in this table means that we skip the word boundary value at every -rd location from the word boundary value list of a given sentence. The pooling is done with the rest of the available word boundaries for these experiments. A higher value of “n” in “Skip-n” indicates an increased use of word boundary information.
column2-4 = c,
hlines,
vline2 = -,
hline3-4 = -dashed,
Stimulus & Skip-2 Skip-3 Skip-4 Skip-4 No Skip
Env. (=1) 74.94 79.10 75.49 80.20 77.45
w2v (=0) 81.1882.78 78.35 80.93 82.00
M-modal 82.69 84.18 80.21 83.47 83.76
@@ III-B3 Region wise Analysis
cells = c,
hlines,
vline2 = -,
hline1 = -0.08em,
hline3,4,5,6 = -dashed,
Brain
Region & Speech
envelope
(=1) Text
word2vec
(=0) Multimodal
(Acoustic
+Semantic)
(=0.5)
Frontal 69.70 74.85 75.65
Central 57.39 62.30 62.47
Parietal 71.09 78.03 78.57
Temporal 77.00 79.33 81.14
Occipital 73.33 79.00 79.98
@@
The electrodes of the -electrode EEG setup are segregated into five regions namely frontal, central, parietal, temporal and occipital. We use the - map to perform this segregation. Only the subset of electrodes belonging to a region are used in training and cross-validation. As shown in Table VI, the temporal region performs significantly better in MM classification in comparison with other regions (Wilcoxon rank-sum test, ).
@@ III-B4 Right Ear Advantage
cells = c,
hlines,
vline2 = -,
hline1,4 = -0.08em,
hline3 = -dashed,
Stimulus & Speech
envelope
(=1) Text
word2vec
(=0) Multimodal
(Acoustic
+Semantic)
(=0.5)
S1 (left ear) 74.14 79.12 83.52
S2 (right ear) 79.55 84.75 87.37
We conducted a set of ablation experiments aimed at identifying which stimulus group (attention to left versus right ear) elicited the strongest performance in the MM classification task. As mentioned in Sec. II-H, the listeners are instructed to attend to stories played in left or right ear. Thus, we split the data into two parts, based on which side of the story the participant was instructed to attend, and chose the trials which positively reflected this instruction in the behavioral data. An initial observation was that the trial selection (discussed in Sec. II-H) retained more trials ( %) when the subjects were asked to attend to the right ear than when they were asked to attend to the left ear (%). The MM task results from the two subject splits are shown in Table VII after balancing the training data of each group to the same number of samples.
We observe that the subject group instructed to listen to the right ear (S2) had significantly improved MM task performance than the subject group (S1) instructed to attend to the left ear (Wilcoxon rank-sum test ). This improvement was consistently seen on all modalities (speech, text and multimodal). We find that this result further aligns with several established studies indicating a right ear advantage in dichotic listening (For example, early work reported in Kimura et al. [41] and more recent studies in Payne et al. [42])
@@ IV Discussion
IV-A Word Boundary Information and Impact on MM Task
With a set of experiments, we have attempted to validate the hypothesis that speech perception in the brain is segmented at the word level. The incorporation of word boundary information yields statistically significant improvements compared to the baseline model, demonstrating the impact of this information in the neural tracking of speech. Both normal and dichotic listening tasks benefit from the use of accurate word information. Further, random choice of word-boundaries do not result in significant improvements.
Our findings suggest that modelling EEG responses and stimuli yield superior performance during natural speech perception. Furthermore, we observed that the relative importance of textual and acoustic content is approximately similar (Wilcoxon signed rank test, ) in the natural listening condition. However, during dichotic listening, the human brain exhibits a prioritization of perceiving the semantics over the acoustic characteristics of the attended speech (Wilcoxon signed rank test, ). The results indicate that word boundary information plays a crucial role in speech perception during the dichotic listening task. Nonetheless, it’s worth noting that recognizing all boundaries during dichotic listening is not as critical compared to the natural speech listening condition, as there is no significant drop in accuracy (Wilcoxon rank-sum test, ) between skip-2 and skip-3 (as well as its subsequent instances)as shown in Table V. This result supports the notion that the human brain prioritizes comprehending the content over phonological aspects of the stimulus during challenging listening conditions, and this comprehension likely occurs at the phrase level or higher.
@@ IV-B Natural Speech: Speech Envelope vs Semantics
Table II displays the outcomes of the MM classification task conducted using two distinct stimulus features: speech envelope and word2vec features. The classification was done at the sentence level using data recorded under normal listening conditions. The results do not show a major difference in performance between the two features (Wilcoxon signed rank test, ). Therefore, it can be concluded that both acoustic and semantic features hold comparable importance in speech perception during normal listening conditions.

### v2

@@ II-F Semantic Feature - Word2vec text embedding
The text transcription of the speech stimulus is provided in the dataset. This study used this text data to obtain features representing the semantics of the speech signal. Semantic vectors for content words were derived using the word2vec algorithm [47].
@@ II-G1 Acoustic Encoding
The EEG sub-network (top row of Figure 2) consists of a series of convolutional layers and LSTM/Transformer layers. The convolutional layers implement -D and -D convolutions with and kernel sizes, respectively. Further, the -D CNN layers also introduce a stride of to further down-sample the feature maps. The speech sub-network (top row in Figure 2) has a 2D convolutional layer of kernel size 16 and stride (1,3). For acoustic encoding, we utilize both the speech and EEG subnetworks in a parallel neural pipeline, as depicted in top half of Figure 2, without any weight sharing.
The word boundary information is incorporated at the equivalent sampling rate (both EEG and audio representations at Hz). The audio and EEG feature maps are average pooled at the word level using the word boundary information. As a result, for a given sentence, the EEG and speech branches generate vector representations sampled at the word level. The LSTM/Transformer layers model the inter-word context from these representations. This layer is included in both the stimulus (speech) and response (EEG) pathways. The output of the LSTM/Transformer layer, of dimension , is used as the embedding for the stimulus/response, denoted as /, respectively.
We propose to optimize the Manhattan distance [49] between the stimulus and response embeddings. The similarity score is computed as,
The similarity score for matched pair () and mismatched pair () are computed. The model, with a dropout of , is trained using a binary cross-entropy loss, with [, ] mapped to [, ] targets.
@@ II-G2 Semantic Encoding
To compare the encoding of text features with speech features, we use a text-EEG MM task. We used the text-EEG sub-network (bottom part of Figure 2) to learn the mapping between the w2v embeddings and the EEG response. The EEG sub-network resembles the acoustic encoding model, while the text sub-network consists of two-layers of LSTM/Transformer architecture with w2v embeddings at word-level as input features. The last hidden state of text sub-network is used as text embedding, denoted as . The similarity scoring component is identical to the acoustic encoding model (Eq. 1).
@@ II-G3 Joint Encoding of Acoustics and Semantics
In addition to exploring the effect of acoustic and semantic features individually, we perform an investigation of the joint training of the STEM3 with both the modalities. The sub-networks are combined by defining a joint-loss function over the two modalities (detailed in Sec. II-I). For a given match segment, the EEG response corresponds to the attended audio/text features, while in the mismatched example, the audio/text features are chosen from a different region (natural speech) or from the stimuli stream played in the other ear (dichotic listening).
To effectively process the Natural Dataset - Large and mitigate the issue of underfitting, the LSTM layer is replaced with a transformer encoder. This modification enhances the model’s capacity to capture complex patterns and dependencies within the larger SparrKULee data. The LSTM architecture is used on Natural Dataset - small and the Dichotic dataset.
@@ II-H Trial selection in Dichotic listening
As part of the dichotic listening task, participants were required to answer multiple-choice questions after each trial to assess their comprehension and, consequently, their attentiveness to the played stories [4]. They were asked questions each from both the stories. The scores obtained were normalized to a range of to . For all the experiments that we report in this work, we include trials in which the participants achieved a comprehension score of % or more for the attended audio, while achieving a score of % or less for the un-attended audio. This choice of scores for attended / un-attended streams ensured that the “matched” / “mismatched” EEG-audio/text pairs had a clear distinction. The selection is elicited in Figure 3. While this selection led to a removal of % of the total trials, the process ensured a reliable label for the training and evaluation.
@@ II-H1 Training and Evaluation Setup
The experiments are run with a batch size of . The models are trained using Adam optimizer [50] with a learning rate of and weight decay parameter of . The models are learned with a binary cross-entropy loss.
Overall, the Natural dataset - Small contains stimulus trials, while the dichotic listening dataset contains trials. A subset of trials from all subjects constituted the training set, while the remaining unseen trials of all subjects served as the test data. In each validation fold, we randomly chose trials to form the test data. This resulted in -fold cross-validation for the natural speech dataset and -fold cross-validation for the cocktail party dataset. The average classification accuracy across validation folds is reported in the results section. In Natural dataset - Large, out of subjects, subjects’ data were kept for evaluation and the remaining were used in training and validation in ratio.
@@ II-I Multimodal Loss Function and Joint training
We explore three choices to combine the modalities,
Here, implies that only the semantic network (text-EEG) is considered, implies that only the acoustic network (speech-EEG) is considered while implies joint modeling (EEG-speech+text). During training, the speech, text and EEG data are loaded and the value of is randomly selected from . In this way, a single model is trained with all three combinations, of text alone (), speech alone () and speech-text multimodal (). During inference, the model can be input with EEG-text samples (), EEG-speech samples () as well EEG-speech+text (). This flexibility helps us understand the role of acoustics and semantics from the same model.
@@ III-A Choice of loss function
For dichotic listening condition, we experiment with different values during inference. As shown in Figure 4, it is observed that semantic encoding in EEG achieves a higher MM performance than the acoustic encoding network. Further, the similarity measure proved to be best performing similarity measure. The best MM classification results are achieved with . All the subsequent experiments use this configuration with similarity measure.
@@ III-B Impact of Acoustic and Semantic Cues
| Speech Envelope | Text Word2vec | Multimodal | ||
| Dataset | Model | (Acoustic) | (Semantic) | (Acoustic + Semantic) |
| Baseline [12] | 65.23 | – | – | |
| Natural Dataset - Small | Borsdorf et al. [21]a | 68.96 | – | – |
| Wang et al. [17]b | – | – | 36.33 | |
| STEM3 | 93.63 | 93.24 | 93.38 | |
| Baseline [42] | 62.00 | – | – | |
| Dichotic Dataset | Borsdorf et al. [21]a | 49.98 | – | – |
| Wang et al. [17]b | – | – | 56.01 | |
| STEM3 | 77.45 | 82.00 | 83.76 | |
| Baseline [12] | 50.05 | – | – | |
| Natural Dataset - Largec | Borsdorf et al. [21]a | 74.77 | – | – |
| Wang et al. [17]b | – | – | 50.33 | |
| STEM3d | 88.70 | 74.37 | 87.91 |
a Envelope with MHA+DC model was selected to maintain uniformity in comparison.
b Envelope + GPT model was selected for consistency; “” does not apply here.
c All results with Natural Dataset - Large were evaluated as described in Section II-A.
d Here, LSTM was replaced with a transformer as discussed in Section II-G.
Table I presents the results of the MM classification accuracy between speech stimuli and their corresponding EEG responses, focusing on different listening conditions. Our STEM3 was evaluated alongside the baseline model [12] [42] and two other models [21] [17] across three datasets. STEM3 demonstrated an average relative improvement of on the Natural Dataset – Small, on the Dichotic dataset, and on the Natural Dataset – Large, with all improvements being statistically significant(Wilcoxon signed-rank test, ) as shown in Table I.
Compared to the MM classification of natural speech under normal listening conditions, the task of auditory attention decoding (AAD) in dichotic listening is more challenging.
For natural speech conditions, both semantic and acoustic features result in similar accuracy (Wilcoxon signed-rank test, ). For dichotic listening, word2vec features result in significant improvement in accuracy compared to speech envelope features (Wilcoxon signed-rank test, ). The scatter plot for prediction probabilites for dichotic listening condition and for a fold in validation is shown in Figure 5.
@@ III-B1 Importance of word boundary information
Word-level segmentation holds significant importance in speech perception during normal listening conditions [22]. To investigate the role of word segmentation in auditory attention decoding, we conducted a comparison of MM classification task performance using the proposed model with and without word boundary information in the dichotic listening condition. The word boundary segmentation is used only for the EEG and acoustic features, while the word2vec features are inherently sampled at word-level.
As seen in Table II, word level segmentation plays an important role in auditory attention decoding. Word boundary information has impact on both semantic features (Wilcoxon rank-sum test, ) and envelope features (Wilcoxon rank-sum test, ). Word boundary information also shows a significant impact on the joint speech-text model.
| Stimulus Feature | Without | With |
| Word Boundary | Word Boundary | |
| Env () | 67.09 | 77.45 |
| w2v () | 77.35 | 82.00 |
| Both () | 76.69 | 83.76 |
@@ III-B2 Importance of accurate word boundaries
We conducted another ablation study to understand the impact of accurate word boundary information. In this setting, the model is fed with random word boundaries with varying granularity. Each sentence is segmented to a fixed number of regions, and their boundaries are chosen at random. The results are reported in Table III. The performance does not significantly improve for speech-EEG (Wilcoxon rank-sum test, ), text-EEG (Wilcoxon rank-sum test, ), or with multimodality (Wilcoxon rank-sum test, ), even as the number of random word boundaries increases. However, the use of accurate word-boundaries is significantly better than those with random word boundaries ().
| Stimulus | 2 | 3 | 4 | 5 | True |
| Feature | Word Boundary | ||||
| Env () | 66.13 | 66.16 | 67.39 | 70.89 | 77.45 |
| w2v () | 78.95 | 76.29 | 78.58 | 78.19 | 82.00 |
| Both () | 79.11 | 79.20 | 79.55 | 79.41 | 83.76 |
@@ III-B3 Region wise Analysis
| Speech | Text | Multimodal | |
| Brain | Envelope | Word2vec | (Acoust. + |
| Region | (Acoust.) | (Sem.) | Sem.) |
| Frontal | 69.70 | 74.85 | 75.65 |
| Central | 57.39 | 62.30 | 62.47 |
| Parietal | 71.09 | 78.03 | 78.57 |
| Temporal | 77.00 | 79.33 | 81.14 |
| Occipital | 73.33 | 79.00 | 79.98 |
@@
The electrodes of the -electrode EEG setup are segregated into five regions namely frontal, central, parietal, temporal and occipital. We use the - map to perform this segregation. Only the subset of electrodes belonging to a region are used in training and cross-validation. As shown in Table IV, the temporal region performs significantly better in MM classification in comparison with other regions (Wilcoxon rank-sum test, ). The performance heat-map (Figure 6), measured for the electrodes belonging to various brain regions, showed a significant preference to the temporal region.
@@ III-B4 Right Ear Advantage
| Speech | Text | Multimodal | |
| Brain | Envelope | Word2vec | (Acoust. + |
| Region | (Acoust.) | (Sem.) | Sem.) |
| S1 (left ear) | 74.14 | 79.12 | 83.52 |
| S2 (right ear) | 79.55 | 84.75 | 87.37 |
We conducted a set of ablation experiments aimed at identifying which stimulus group (attention to left versus right ear) elicited the strongest performance in the MM classification task for dichotic listening task. As mentioned in Section II-H, the listeners are instructed to attend to stories played in left or right ear. Thus, we split the data into two parts, based on the story the participant was instructed to attend, and chose the trials which positively reflected this instruction in the behavioral data. An initial observation was that the trial selection (discussed in Section II-H) retained more trials ( %) when the subjects were asked to attend to the right ear than when they were asked to attend to the left ear (%). The MM task results from the two subject splits are shown in Table V, after balancing the training data of each group to the same number of samples.
We observed that the subject group instructed to listen to the right ear (S2) had significantly improved MM task performance than the subject group (S1) instructed to attend to the left ear (Wilcoxon rank-sum test ). This improvement was consistently seen on all modalities (speech, text and multimodal).
@@ IV Discussion
IV-A Word Boundary Information
With a set of experiments, we have attempted to validate the hypothesis that speech perception in the brain is segmented at the word level. The incorporation of word boundary information yields statistically significant improvements compared to the baseline model, demonstrating the impact of this information in the neural tracking of speech. Both normal and dichotic listening tasks benefit from the use of accurate word information. Further, random choice of word-boundaries does not have any significant impact on the task. This result supports the notion that the human brain prioritizes comprehending the content over phonological aspects of the stimulus during challenging listening conditions, and this comprehension likely occurs at the word level.
@@ IV-B Natural Speech: Speech Envelope vs Semantics
Table I displays the outcomes of the MM classification task conducted using two distinct stimulus features: speech envelope and word2vec features. Our findings suggest that modelling EEG responses and stimuli yield superior performance during natural speech perception. Furthermore, we observed that the relative importance of textual and acoustic content is approximately similar (Wilcoxon signed-rank test, ) in the natural listening condition (small). Therefore, it can be concluded that both acoustic and semantic features hold comparable importance in speech perception during normal listening conditions. In natural listening condition (large), we find the acoustic features to be significantly better than textual features (Wilcoxon signed-rank test, ).